Circular No. 7186

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

MARS T. Encrenaz and E. Lellouch, DESPA, Paris Observatory; G. Paubert,

Institut de Radio-Astronomie Millimetrique (IRAM), Granada; and S. Gulkis, Jet Propulsion Laboratory, report: “On Apr. 25–28, we detected a millimeter rotational transition of NO at frequency 250.436848 GHz in the atmosphere of Mars, using the IRAM 30-m antenna at Pico Veleta, Spain. This is, to our knowledge, the first ground-based detection of a nitrogenbearing species in the Martian atmosphere. The observed line position is found to be within 30 kHz of the position predicted by the Submillimeter, millimeter and microwave spectra catalog (Pickett et al. 1992, JPL Tech. Publ. 80-23, revision 3). The line is observed in emission, and the line width corresponds to a Doppler-broadened line, indicating that the line is formed in Mars’ upper atmosphere. The line intensity appears, at first order, consistent with the synthetic calculations of Encrenaz et al. (1995, Planet. Space Sci. 43, 1485), based on the NO vertical distribution inferred from the Viking measurements.”

SUPERNOVA 1999bv IN MCG +10-25-14 G. J. Hill, McDonald Observatory, University of Texas, reports on

behalf of the Hobby-Eberly Telescope (HET) Low Resolution Spectrograph (LRS) commissioning team (which also includes M. P. Bergmann, D. A. Howell, J. C. Wheeler, P. J. MacQueen, J. A. Booth, M. T. Adams, M. Eracleus, C. Tejada, F. Cobos, F. Garfias, J. Cuervo, H. Nicklas, W. Mitsch, G. W. Wesley, P. S. Odoms, G. Barczac, L. W. Ramsey, T. G. Barnes III, B. Roman, C. Nance, G. Hill, T. Worthington): “A low-resolution spectrogram (range 400–1000 nm) of SN 1999bv was obtained on Apr. 21.37 UT with the Marcario LRS, during first light commissioning of the instrument. The broad emission features seen by Jha et al. (IAUC 7150) at 667, 600, 475, and 402 nm are confirmed, and we note that the broad emission (FWHM 10 000 km/s) at 667 nm is consistent with slightly blueshifted Hα emission in the rest frame of the supernova. Such emission was also seen in SN 1993J, and is suspected to be an indication of transition, or hybrid type-II/type-Ib/c events, sometimes referred to as type IIb.”

Corrigendum. On IAUC 7150, for MGC read MCG

1999 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green